what is the general solution to sin(x/2)=(root3/2)
Answer by venugopalramana(3286)

what is the general solution to sin(x/2)=(root3/2)
X/2=N*PI+(-1)^N*(PI/3)
X=2N*PI+(-1)^N*(2PI/3)